Your answer is incorrect. Try again. Monty Inc., a manufacturer of steel school lockers, plans to purchase a new punch press for use in its manufacturing process. After contacting the appropriate vendors, the purchasing department received differing terms and options from each vendor. The Engineering Department has determined that each vendor’s punch press is substantially identical and each has a useful life of 20 years. In addition, Engineering has estimated that required year-end maintenance costs will be $900 per year for the first 5 years, $1,900 per year for the next 10 years, and $2,900 per year for the last 5 years. Following is each vendor’s sales package. Vendor A: $57,710 cash at time of delivery and 10 year-end payments of $17,430 each. Vendor A offers all its customers the right to purchase at the time of sale a separate 20-year maintenance service contract, under which Vendor A will perform all year-end maintenance at a one-time initial cost of $10,510. Vendor B: Forty semiannual payments of $10,100 each, with the first installment due upon delivery. Vendor B will perform all year-end maintenance for the next 20 years at no extra charge. Vendor C: Full cash price of $158,200 will be due upon delivery. Assuming that both Vendors A and B will be able to perform the required year-end maintenance, that Monty’s cost of funds is 10%, and the machine will be purchased on January 1, compute the following: Click here to view factor tables The present value of the cash flows for vendor A. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.) The present value of the cash outflows for this option is--- The present value of the cash flows for vendor B. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.) The present value of the cash outflows for this option is----- The present value of the cash flows for vendor C. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.) The present value of the cash outflows for this option is ------- From which vendor should the press be purchased? The press should be purchased from------
Answer #1

Vendor A

Year Cash outflow for press Cash outflow for repairs Total Cash outflow Present value factors @10% Present value of cash flows
0 57710 10510 68220 1 68220
1 17430 17430 0.90909 15845.44
2 17430 17430 0.82645 14405.02
3 17430 17430 0.75131 13095.33
4 17430 17430 0.68301 11904.86
5 17430 17430 0.62092 10822.64
6 17430 17430 0.56447 9838.712
7 17430 17430 0.51316 8944.379
8 17430 17430 0.46651 8131.269
9 17430 17430 0.4241 7392.063
10 17430 17430 0.38554 6719.962
175320

The present value of the cash outflows for this option is $175,320.

Vendor B

Present value of annuity = A\frac{1-(1+r)^{-n}}{r} = 10100\frac{1-1.05^{-40}}{.05} = 10100 * 17.15909 = 173307

The present value of the cash outflows for this option is $173,307

Vendor C

Year Cash outflow for press Cash outflow for repairs Total Cash outflow Present value factors @10% Present value of cash flows
0 158200 0 158200 1 158200
1 900 900 0.90909 818.181
2 900 900 0.82645 743.805
3 900 900 0.75131 676.179
4 900 900 0.68301 614.709
5 900 900 0.62092 558.828
6 1900 1900 0.56447 1072.493
7 1900 1900 0.51316 975.004
8 1900 1900 0.46651 886.369
9 1900 1900 0.4241 805.79
10 1900 1900 0.38554 732.526
11 1900 1900 0.35049 665.931
12 1900 1900 0.31863 605.397
13 1900 1900 0.28966 550.354
14 1900 1900 0.26333 500.327
15 1900 1900 0.23939 454.841
16 2900 2900 0.21763 631.127
17 2900 2900 0.19784 573.736
18 2900 2900 0.17986 521.594
19 2900 2900 0.16351 474.179
20 2900 2900 0.14864 431.056
171492

The present value of the cash outflows for this option is $171,492.

Thus, the press should be purchased from vendor C

Kindly Up vote